OpenStudy (anonymous):

write a word problem that can be solved using the equation 5x=3x+20

OpenStudy (katrinakaif):

Have you tried to create one?

hero (hero):

John has 5 times a certain number of cookies. Jerry has 3 times a certain number of cookies plus twenty more. They both have the same amount of cookies. How many cookies does each person have?

hero (hero):

Let x = a certain number of cookies 5x = 3x + 20 5x - 3x = 20 2x = 20 x = 20/2 x = 10 John: 5x = 5(10) = 50 Jerry: 3x + 20 = 3(10) + 20 = 30 + 20 = 50 They both have 50 cookies a piece

hero (hero):

At least one of them will get sick from eating all their cookies.

hero (hero):

I guess you didn't like my word problem.

OpenStudy (katrinakaif):

Lol